# Combinatorial flats and Schubert varieties of subspace arrangements

### v2

@@ 4. Schubert varieties of subspace arrangements
Let $\mathbf{s}\leq\mathbf{n}$ be a multiset represented by $S\subset\widetilde{E}$. From Section 2.5, we learn
@@
We establish 1.3 via a series of lemmas characterizing the strata of $Y_{V}$. Throughout this section, $V$ is assumed to be polymatroid general.
@@
Otherwise, suppose $N>1$. If $\operatorname{rk}(\mathbf{s})=d$, then the dimension formula implies $Y_{V}\cap O_{\mathbf{s}}$ is nonempty if and only if $\mathbf{s}=\mathbf{n}$. Otherwise, if $\operatorname{rk}\mathbf{s}<d$, then there is $i\in E$ such that $\overline{s}_{i}<n_{i}$. The projection $Y_{V}\to Y_{\pi_{E\setminus i}(V)}$ sends $Y_{V}\cap O_{\mathbf{s}}$ into $Y_{\pi_{E\setminus i}(V)}\cap O_{\pi_{E\setminus i}(\mathbf{s})}$. By 3.9, $\pi_{E\setminus i}(\mathbf{s})$ is not a combinatorial flat of $\pi_{E\setminus i}(V)$, so $Y_{\pi_{E\setminus i}(V)}\cap O_{\pi_{E\setminus i}(\mathbf{s})}$ is empty by induction on $N$. This means $Y_{V}\cap O_{\mathbf{s}}$ is empty too. ∎
@@
We complete the proof for proper combinatorial flats by inducting on $N$. First consider $N=1$, so that $Y_{V}\subset\mathbb{P}^{n_{1}}$. By 4.10, $Y_{V}\cap\mathbb{P}^{n_{1}-1}\subset\mathbb{P}^{\dim V-1}$. Since $Y_{V}\cap\mathbb{P}^{n_{1}-1}$ is a divisor on $Y_{V}$, we must have $Y_{V}\cap\mathbb{P}^{n_{1}-1}=\mathbb{P}^{\dim V-1}$. Since $V$ is polymatroid general, $O_{i}\subset\mathbb{P}^{\dim V-1}$ is a single $V$-orbit, and $\overline{Y_{V}\cap O_{i}}=\mathbb{P}^{i}$ as desired.
@@
The dimension of the left-hand set is at least $\operatorname{rk}(\mathbf{s})$ because $V$ is polymatroid general; the right-hand set is connected of dimension $\operatorname{rk}(\mathbf{s})$ by 3.2. Hence, the two sides are equal.
@@
We induct on $\dim V$. If $\dim V=1$, let $\mathbf{s}$ be its minimal combinatorial flat. Since $V$ is just a line, one sees that $O_{\mathbf{s}}\subset Y_{V}$ using the explicit description of the Bell polynomials in Section 4.1.
@@
Otherwise, suppose $\dim V>1$. For a general hyperplane $H$ in $\mathbb{K}^{\mathbf{n}}$, $Y_{V\cap H}\cap O_{\mathbf{s}}$ is nonempty for every combinatorial flat $\mathbf{s}$ of corank greater than 1 by the induction hypothesis and 3.10. By the same argument, if $F=\{i:s_{i}=n_{i}\}$ is nonempty, then $Y_{V\cap H}\cap O_{\mathbf{s}}$ is nonempty for $H\subset\mathbb{K}^{\mathbf{n}}$ the preimage of a general hyperplane in $\mathbb{K}^{\pi_{F}(\mathbf{n})}$.
@@
We have now reduced to the case where $\mathbf{s}$ is a combinatorial flat of rank $d-1$ with $\mathbf{s}^{\mathrm{geo}}=\mathbf{0}$. By 3.4, this means that $\mathbf{s}$ is an independent multiset of cardinality $d-1$, so there is $i\in E$ such that $\mathbf{s}+\mathbf{e}_{i}$ is a basis of $P$. Let $H=\mathbb{P}^{\mathbf{n}-\mathbf{e}_{i}}$. Every irreducible component of $Y_{V}\cap H$ is $(d-1)$-dimensional, so by [EH16, Theorem 1.26],
$[\mathbb{P}^{\mathbf{n}-\mathbf{e}_{i}}]\cdot[Y_{V}]=\sum_{C}m_{C}[C],$
@@
where $C$ runs over all irreducible components of $Y_{V}\cap H$. By 4.10, $Y_{V}\cap H\subset\cup_{\mathbf{s}^{\prime}}Y_{V}\cap\mathbb{P}^{\mathbf{s}^{\prime}}$, where the union runs over all corank 1 combinatorial flats $\mathbf{s}^{\prime}\leq\mathbf{n}-\mathbf{e}_{i}$. By 4.11, each irreducible component $C$ of $Y_{V}\cap H$ is among those of $\cup_{\mathbf{s}^{\prime}}Y_{V}\cap\mathbb{P}^{\mathbf{s}^{\prime}}$, so
$[\mathbb{P}^{\mathbf{n}-\mathbf{e}_{i}}]\cdot[Y_{V}]=\sum_{\mathbf{s}^{\prime}}m_{\mathbf{s}^{\prime}}[Y_{V}\cap\mathbb{P}^{\mathbf{s}^{\prime}}],$
with $m_{\mathbf{s}^{\prime}}=0$ if $\dim Y_{V}\cap\mathbb{P}^{\mathbf{s}^{\prime}}<d-1$. By 4.11 and 4.6, we may further expand as
$[\mathbb{P}^{\mathbf{n}-\mathbf{e}_{i}}]\cdot[Y_{V}]=\sum_{\mathbf{s}^{\prime}}m_{\mathbf{s}^{\prime}}\sum_{\mathbf{b}^{\prime}\text{ basis of $\mathbf{s}^{\prime}$}}c_{\mathbf{s}^{\prime},\mathbf{b}^{\prime}}[\mathbb{P}^{\mathbf{b}^{\prime}}].$
On the other hand, by 4.6,
$[\mathbb{P}^{\mathbf{n}-\mathbf{e}_{i}}]\cdot[Y_{V}]=\sum_{\begin{subarray}{c}\mathbf{b}\text{ basis of $P$}\\ b_{i}>0\end{subarray}}c_{\mathbf{b}}[\mathbb{P}^{\mathbf{b}-\mathbf{e}_{i}}],$
@@
with all coefficients positive. Since $\mathbf{s}$ is both an independent multiset and a combinatorial flat of $P$, $[\mathbb{P}^{\mathbf{s}}]$ appears in both expansions, and $0<c_{\mathbf{s}}=m_{\mathbf{s}}c_{\mathbf{s},\mathbf{s}}$. This implies $m_{\mathbf{s}}\neq 0$, so $\dim Y_{V}\cap\mathbb{P}^{\mathbf{s}}=d-1$, hence $Y_{V}\cap O_{\mathbf{s}}\neq\emptyset$. ∎
Proof of 1.3.
@@
Characterization of when $Y_{V}$ intersects $O_{\mathbf{s}}$ is 4.10 and 4.12. 1.3i follows from 4.11, and 1.3iii is obtained by combining 4.11 and 4.6.
@@
It remains to prove 1.3ii, which states that if $\mathbf{s},\mathbf{s}^{\prime}\in\mathcal{L}_{P,\mathbf{n}}$, then $\overline{Y_{V}\cap O_{\mathbf{s}}}\supset Y_{V}\cap O_{\mathbf{s}^{\prime}}$ if and only if $\mathbf{s}\geq\mathbf{s}^{\prime}$.
@@
Fix $\mathbf{s}\in\mathcal{L}_{P,\mathbf{n}}$. By 3.9, $\mathbf{s}^{\prime}<\mathbf{s}$ is a combinatorial flat of $V$ if and only if $\mathbf{s}_{i}^{\prime}\leq\mathbf{s}_{i}$ for all $i$ and $(\mathbf{s}_{j}^{\prime})_{j:s_{j}=n_{j}}$ is a combinatorial flat of $V^{\prime}:=\pi_{\{j:s_{j}=n_{j}\}}(V)$. Such multisets $\mathbf{s}^{\prime}$ are precisely those that index the nonempty strata of $\overline{Y_{V}\cap O_{\mathbf{s}}}$, since $\overline{Y_{V}\cap O_{\mathbf{s}}}=Y_{V^{\prime}}\times\prod_{s_{i}<n_{i}}\mathbb{P}^{s_{i}}$ by 4.11. 4.11 also implies $O_{\mathbf{s}^{\prime}}\cap(Y_{V^{\prime}}\times\prod_{s_{i}<n_{i}}\mathbb{P}^{s_{i}})=O_{\mathbf{s}^{\prime}}\cap Y_{V}$ whenever both sides are nonempty, which completes the proof. ∎
@@
Theorems 1.3 describes properties of Schubert varieties of subspace arrangements, extending well-known results for Schubert varieties of hyperplane arrangements from [AB16], which correspond to the case $\mathbf{n}=(1,\ldots,1)$. One further such property follows: when $\mathbf{n}=(1,\ldots,1)$, it is well-known that changing $V$ by rescaling each coordinate of $\mathbb{K}^{\mathbf{n}}=\mathbb{K}^{N}$ does not change the isomorphism class of $Y_{V}$. For general $\mathbf{n}$, the isomorphism class of $Y_{V}$ is unchanged if we perform a weighted rescaling of each factor $\mathbb{K}^{n_{i}}$, as described in Section 4.2.
@@
Assume that $V$ realizes a simple polymatroid. The subspaces $V_{ij}=V\cap\operatorname{St}_{j\mathbf{e}_{i}}$ for $1\leq i\leq N$ and $1\leq j\leq n_{i}$ comprise a collection of $N$ flags in $V$. The combinatorial flats record intersection pattern of these flags, as $\operatorname{rk}(\mathbf{s})=\operatorname{codim}_{V}V_{1,s_{1}}\cap V_{2,s_{2}}\cap\cdots\cap V_{N,s_{N}}$. The “polymatroid general” condition on $V$ guarantees that these flags are in general position with respect to one another (see also 3.7).
@@
The data $V\subset\mathbb{K}^{\mathbf{n}}$ realizes the nonsimple polymatroid of 3.18, whose simplification is the boolean matroid on two elements. However, we will see that $Y_{V}$ is not isomorphic to the Schubert variety $(\mathbb{P}^{1})^{2}$ of the simplification. Consider the linear paramitrization $(u,v)\mapsto(u,v,u+v,u-v)$ of $V$, and the corresponding paramatrization of $\iota(V)\subset\mathbb{P}^{2}\times\mathbb{P}^{2}$:
@@
Recall that we label the coordinates on the right hand side as $[b_{10}:b_{11}:b_{12}]\times[b_{20}:b_{21}:b_{22}]$. The relations between the coordinates $b_{11},b_{12},b_{21},b_{22}$ on the right hand side are generated by the inhomogenious equations
@@
In this section, we establish 1.7, which gives a formula for the cohomology ring of $Y_{V}$. Throughout, we assume that $V$ is polymatroid general. For an independent multiset $\mathbf{b}$, let $c_{\mathbf{b}}>0$ be the coefficient of $[\mathbb{P}^{\mathbf{b}}]$ in $[Y_{V}\cap\mathbb{P}^{\overline{\mathbf{b}}}]$, as in the statement of 1.3.
@@
Let $\kappa:Y_{V}\to\mathbb{P}^{\mathbf{n}}$ denote the inclusion. By 1.3iii, the classes $\{[Y_{V}\cap\mathbb{P}^{\mathbf{s}}]:\mathbf{s}\in\mathcal{L}_{P,\mathbf{n}}\}$ are linearly independent in $A_{*}\mathbb{P}^{\mathbf{n}}$, so the pushforward $A_{*}Y_{V}\to A_{*}\mathbb{P}^{\mathbf{n}}$ is injective.
@@
Let us verify that the remaining generators of $I$ are in $\ker(\kappa^{*})$. By the Universal Coefficient Theorem, the pullback of $\alpha\in H^{k}(\mathbb{P}^{\mathbf{n}};\mathbb{Q})$ is zero if and only if its cap product with any element of $H_{k}(Y_{V};\mathbb{Q})$ is zero. By the projection formula and the injectivity of $\kappa_{*}$, this is equivalent to $\alpha\smallfrown[\overline{Y_{V}\cap O_{\mathbf{s}}}]=0$ for all rank $k$ combinatorial flats $\mathbf{s}$ of $P$. By 1.3iii,
@@
Proof of 1.7.
Let $S=\mathbb{Q}[y_{\mathbf{s}}:\text{$\mathbf{s}\in\mathcal{L}_{P,\mathbf{n}}$}]$, graded by $\deg y_{\mathbf{s}}=\operatorname{rk}_{P}(\mathbf{s})$. Using the notation of 4.16, define a map
@@
where $\mathbf{b}$ is any basis of $\mathbf{s}$. The image of $y_{\mathbf{s}}$ does not depend on $\mathbf{b}$ by 4.16.
@@
In general, our proof of 1.1iv (Section 3.4) relies on high-powered results of [BHM+20]. However, if $P$ is realized by $V\subset\mathbb{K}^{\mathbf{n}}$ with $\mathbb{K}$ of characteristic 0, then our work gives an independent proof of top-heaviness for $\mathcal{L}_{P}$ via arguments similar to those contained in [HW17, BE09].
@@
are injective when $i\leq d/2$. These maps preserve the submodule $H^{*}(Y_{V})\subset I\!H^{*}(Y_{V})$, so they restrict to injections $H^{i}(Y_{V})\to H^{d-i}(Y_{V})$. 1.7 states that $\dim H^{k}(Y_{V})=|\mathcal{L}_{P}^{k}|$, so taking dimensions yields top-heaviness for $\mathcal{L}_{P}$.

### v3

@@ 4. Schubert varieties of subspace arrangements
Let $\mathbf{s}\leq\mathbf{n}$ be a multiset represented by $S\subset\widetilde{E}$. From Section˜2.5, we learn
@@
We establish ˜1.3 via a series of lemmas characterizing the strata of $Y_{V}$. Throughout this section, $V$ is assumed to be polymatroid general.
@@
Otherwise, suppose $N>1$. If $\operatorname{rk}(\mathbf{s})=d$, then the dimension formula implies $Y_{V}\cap O_{\mathbf{s}}$ is nonempty if and only if $\mathbf{s}=\mathbf{n}$. Otherwise, if $\operatorname{rk}\mathbf{s}<d$, then there is $i\in E$ such that $\overline{s}_{i}<n_{i}$. The projection $Y_{V}\to Y_{\pi_{E\setminus i}(V)}$ sends $Y_{V}\cap O_{\mathbf{s}}$ into $Y_{\pi_{E\setminus i}(V)}\cap O_{\pi_{E\setminus i}(\mathbf{s})}$. By ˜3.9, $\pi_{E\setminus i}(\mathbf{s})$ is not a combinatorial flat of $\pi_{E\setminus i}(V)$, so $Y_{\pi_{E\setminus i}(V)}\cap O_{\pi_{E\setminus i}(\mathbf{s})}$ is empty by induction on $N$. This means $Y_{V}\cap O_{\mathbf{s}}$ is empty too. ∎
@@
We complete the proof for proper combinatorial flats by inducting on $N$. First consider $N=1$, so that $Y_{V}\subset\mathbb{P}^{n_{1}}$. By ˜4.10, $Y_{V}\cap\mathbb{P}^{n_{1}-1}\subset\mathbb{P}^{\dim V-1}$. Since $Y_{V}\cap\mathbb{P}^{n_{1}-1}$ is a divisor on $Y_{V}$, we must have $Y_{V}\cap\mathbb{P}^{n_{1}-1}=\mathbb{P}^{\dim V-1}$. Since $V$ is polymatroid general, $O_{i}\subset\mathbb{P}^{\dim V-1}$ is a single $V$-orbit, and $\overline{Y_{V}\cap O_{i}}=\mathbb{P}^{i}$ as desired.
@@
The dimension of the left-hand set is at least $\operatorname{rk}(\mathbf{s})$ because $V$ is polymatroid general; the right-hand set is connected of dimension $\operatorname{rk}(\mathbf{s})$ by ˜3.2. Hence, the two sides are equal.
@@
We induct on $\dim V$. If $\dim V=1$, let $\mathbf{s}$ be its minimal combinatorial flat. Since $V$ is just a line, one sees that $O_{\mathbf{s}}\subset Y_{V}$ using the explicit description of the Bell polynomials in Section˜4.1.
@@
Otherwise, suppose $\dim V>1$. For a general hyperplane $H$ in $\mathbb{K}^{\mathbf{n}}$, $Y_{V\cap H}\cap O_{\mathbf{s}}$ is nonempty for every combinatorial flat $\mathbf{s}$ of corank greater than 1 by the induction hypothesis and ˜3.10. By the same argument, if $F=\{i:s_{i}=n_{i}\}$ is nonempty, then $Y_{V\cap H}\cap O_{\mathbf{s}}$ is nonempty for $H\subset\mathbb{K}^{\mathbf{n}}$ the preimage of a general hyperplane in $\mathbb{K}^{\pi_{F}(\mathbf{n})}$.
@@
We have now reduced to the case where $\mathbf{s}$ is a combinatorial flat of rank d1d-1 with $\mathbf{s}^{\mathrm{geo}}=\mathbf{0}$. By ˜3.4, this means that $\mathbf{s}$ is an independent multiset of cardinality d1d-1, so there is $i\in E$ such that $\mathbf{s}+\mathbf{e}_{i}$ is a basis of $P$. Let $H=\mathbb{P}^{\mathbf{n}-\mathbf{e}_{i}}$. Every irreducible component of $Y_{V}\cap H$ is (d1)(d-1)-dimensional, so by [EH16, Theorem 1.26],
$\mathbb{P}^{\mathbf{n}-\mathbf{e}_{i}}]\cdot[Y_{V}]=\sum_{C}m_{C}[C],$
@@
where $C$ runs over all irreducible components of $Y_{V}\cap H$. By ˜4.10, $Y_{V}\cap H\subset\cup_{\mathbf{s}^{\prime}}Y_{V}\cap\mathbb{P}^{\mathbf{s}^{\prime}}$, where the union runs over all corank 1 combinatorial flats $\mathbf{s}^{\prime}\leq\mathbf{n}-\mathbf{e}_{i}$. By ˜4.11, each irreducible component $C$ of $Y_{V}\cap H$ is among those of $\cup_{\mathbf{s}^{\prime}}Y_{V}\cap\mathbb{P}^{\mathbf{s}^{\prime}}$, so
$\mathbb{P}^{\mathbf{n}-\mathbf{e}_{i}}]\cdot[Y_{V}]=\sum_{\mathbf{s}^{\prime}}m_{\mathbf{s}^{\prime}}[Y_{V}\cap\mathbb{P}^{\mathbf{s}^{\prime}}],$
with $m_{\mathbf{s}^{\prime}}=0$ if $\dim Y_{V}\cap\mathbb{P}^{\mathbf{s}^{\prime}}<d-1$. By ˜4.11 and ˜4.6, we may further expand as
[…]
On the other hand, by ˜4.6,
[…]
@@
with all coefficients positive. Since $\mathbf{s}$ is both an independent multiset and a combinatorial flat of $P$, $\mathbb{P}^{\mathbf{s}}]$ appears in both expansions, and $0<c_{\mathbf{s}}=m_{\mathbf{s}}c_{\mathbf{s},\mathbf{s}}$. This implies $m_{\mathbf{s}}\neq 0$, so $\dim Y_{V}\cap\mathbb{P}^{\mathbf{s}}=d-1$, hence $Y_{V}\cap O_{\mathbf{s}}\neq\emptyset$. ∎
Proof of ˜1.3.
@@
Characterization of when $Y_{V}$ intersects $O_{\mathbf{s}}$ is ˜4.10 and ˜4.12. ˜1.3˜i follows from ˜4.11, and ˜1.3˜iii is obtained by combining ˜4.11 and ˜4.6.
@@
It remains to prove ˜1.3˜$i$, which states that if $\mathbf{s},\mathbf{s}^{\prime}\in\mathcal{L}_{P,\mathbf{n}}$, then $\overline{Y_{V}\cap O_{\mathbf{s}}}\supset Y_{V}\cap O_{\mathbf{s}^{\prime}}$ if and only if $\mathbf{s}\geq\mathbf{s}^{\prime}$.
@@
Fix $\mathbf{s}\in\mathcal{L}_{P,\mathbf{n}}$. By ˜3.9, $\mathbf{s}^{\prime}<\mathbf{s}$ is a combinatorial flat of $V$ if and only if $\mathbf{s}_{i}^{\prime}\leq\mathbf{s}_{i}$ for all $i$ and $(\mathbf{s}_{j}^{\prime})_{j:s_{j}=n_{j}}$ is a combinatorial flat of $V^{\prime}:=\pi_{\{j:s_{j}=n_{j}\}}(V)$. Such multisets $\mathbf{s}^{\prime}$ are precisely those that index the nonempty strata of $\overline{Y_{V}\cap O_{\mathbf{s}}}$, since $\overline{Y_{V}\cap O_{\mathbf{s}}}=Y_{V^{\prime}}\times\prod_{s_{i}<n_{i}}\mathbb{P}^{s_{i}}$ by ˜4.11. ˜4.11 also implies $O_{\mathbf{s}^{\prime}}\cap(Y_{V^{\prime}}\times\prod_{s_{i}<n_{i}}\mathbb{P}^{s_{i}})=O_{\mathbf{s}^{\prime}}\cap Y_{V}$ whenever both sides are nonempty, which completes the proof. ∎
@@
Theorems 1.3 describes properties of Schubert varieties of subspace arrangements, extending well-known results for Schubert varieties of hyperplane arrangements from [AB16], which correspond to the case $\mathbf{n}=(1,\ldots,1)$. One further such property follows: when $\mathbf{n}=(1,\ldots,1)$, it is well-known that changing $V$ by rescaling each coordinate of $\mathbb{K}^{\mathbf{n}}=\mathbb{K}^{N}$ does not change the isomorphism class of $Y_{V}$. For general $\mathbf{n}$, the isomorphism class of $Y_{V}$ is unchanged if we perform a weighted rescaling of each factor $\mathbb{K}^{n_{i}}$, as described in Section˜4.2.
@@
Assume that $V$ realizes a simple polymatroid. The subspaces $V_{ij}=V\cap\operatorname{St}_{j\mathbf{e}_{i}}$ for $1\leq i\leq N$ and $1\leq j\leq n_{i}$ comprise a collection of $N$ flags in $V$. The combinatorial flats record intersection pattern of these flags, as $\operatorname{rk}(\mathbf{s})=\operatorname{codim}_{V}V_{1,s_{1}}\cap V_{2,s_{2}}\cap\cdots\cap V_{N,s_{N}}$. The “polymatroid general” condition on $V$ guarantees that these flags are in general position with respect to one another (see also ˜3.7).
@@
The data $V\subset\mathbb{K}^{\mathbf{n}}$ realizes the nonsimple polymatroid of ˜3.18, whose simplification is the boolean matroid on two elements. However, we will see that $Y_{V}$ is not isomorphic to the Schubert variety $(\mathbb{P}^{1})^{2}$ of the simplification. Consider the linear paramitrization $(u,v)\mapsto(u,v,u+v,u-v)$ of $V$, and the corresponding paramatrization of $\iota(V)\subset\mathbb{P}^{2}\times\mathbb{P}^{2}$:
@@
Recall that we label the coordinates on the right hand side as $b_{10}:b_{11}:b_{12}]\times[b_{20}:b_{21}:b_{22}]$. The relations between the coordinates $b_{11},b_{12},b_{21},b_{22}$ on the right hand side are generated by the inhomogenious equations
@@
In this section, we establish ˜1.7, which gives a formula for the cohomology ring of $Y_{V}$. Throughout, we assume that $V$ is polymatroid general. For an independent multiset $\mathbf{b}$, let $c_{\mathbf{b}}>0$ be the coefficient of $\mathbb{P}^{\mathbf{b}}]$ in $Y_{V}\cap\mathbb{P}^{\overline{\mathbf{b}}}]$, as in the statement of ˜1.3.
@@
Let $\kappa:Y_{V}\to\mathbb{P}^{\mathbf{n}}$ denote the inclusion. By ˜1.3˜iii, the classes $\{[Y_{V}\cap\mathbb{P}^{\mathbf{s}}]:\mathbf{s}\in\mathcal{L}_{P,\mathbf{n}}\}$ are linearly independent in $A_{*}\mathbb{P}^{\mathbf{n}}$, so the pushforward $A_{*}Y_{V}\to A_{*}\mathbb{P}^{\mathbf{n}}$ is injective.
@@
Let us verify that the remaining generators of $I$ are in $\ker(\kappa^{*})$. By the Universal Coefficient Theorem, the pullback of $\alpha\in H^{k}(\mathbb{P}^{\mathbf{n}};\mathbb{Q})$ is zero if and only if its cap product with any element of $H_{k}(Y_{V};\mathbb{Q})$ is zero. By the projection formula and the injectivity of $\kappa_{*}$, this is equivalent to $\alpha\smallfrown[\overline{Y_{V}\cap O_{\mathbf{s}}}]=0$ for all rank $k$ combinatorial flats $\mathbf{s}$ of $P$. By ˜1.3˜iii,
@@
Proof of ˜1.7.
Let $S=\mathbb{Q}[y_{\mathbf{s}}:\text{$\mathbf{s}\in\mathcal{L}_{P,\mathbf{n}}$}]$, graded by $\deg y_{\mathbf{s}}=\operatorname{rk}_{P}(\mathbf{s})$. Using the notation of ˜4.16, define a map
@@
where $\mathbf{b}$ is any basis of $\mathbf{s}$. The image of $y_{\mathbf{s}}$ does not depend on $\mathbf{b}$ by ˜4.16.
@@
In general, our proof of ˜1.1˜iv (Section˜3.4) relies on high-powered results of [BHM+20]. However, if $P$ is realized by $V\subset\mathbb{K}^{\mathbf{n}}$ with $\mathbb{K}$ of characteristic 0, then our work gives an independent proof of top-heaviness for $\mathcal{L}_{P}$ via arguments similar to those contained in [HW17, BE09].
@@
are injective when $i\leq d/2$. These maps preserve the submodule $H^{*}(Y_{V})\subset I\!H^{*}(Y_{V})$, so they restrict to injections $H^{i}(Y_{V})\to H^{d-i}(Y_{V})$. ˜1.7 states that $\dim H^{k}(Y_{V})=|\mathcal{L}_{P}^{k}|$, so taking dimensions yields top-heaviness for $\mathcal{L}_{P}$.